# Six-CD: Benchmarking Concept Removals for Benign Text-to-image Diffusion Models

### v1

@@ Abstract
Text-to-image (T2I) diffusion models have shown exceptional capabilities in generating images that closely correspond to textual prompts. However, the advancement of T2I diffusion models presents significant risks, as the models could be exploited for malicious purposes, such as generating images with violence or nudity, or creating unauthorized portraits of public figures in inappropriate contexts. To mitigate these risks, concept removal methods have been proposed. These methods aim to modify diffusion models to prevent the generation of malicious and unwanted concepts. Despite these efforts, existing research faces several challenges: (1) a lack of consistent comparisons on a comprehensive dataset, (2) ineffective prompts in harmful and nudity concepts, (3) overlooked evaluation of the ability to generate the benign part within prompts containing malicious concepts. To address these gaps, we propose to benchmark the concept removal methods by introducing a new dataset, Six-CD, along with a novel evaluation metric. In this benchmark, we conduct a thorough evaluation of concept removals, with the experimental observations and discussions offering valuable insights in the field. The dataset and code is available at github.com/Artanisax/Six-CD.
@@ 1 Introduction
Strategies have been proposed to mitigate the malicious generation. For example, in the data collection stage, Stable Diffusion (SD) [6] and Adobe [16] filter out the Not-Safe-For-Work (NSFW) contents in training data. After image generation, SD uses a detector to ensure the images are benign, removing unwanted ones. However, training data filtering requires retraining and will be inefficient if the model has already been trained on various data. For open-sourced models like SD [6], the detector can be easily disabled by modifying the source code. Therefore, concept removal techniques have been proposed to edit a trained model’s parameters to prevent generating malicious concepts. For instance, methods are proposed to fine-tune the models to replace unwanted concepts with benign alternatives [17, 18, 19]. Besides, to accelerate fine-tuning, some methods fine-tune the linear components (e.g. the weight matrix of cross attention) using the closed-form optimal solution [20, 21, 22]. Alternatively, there are methods modifying the output during inference to bypass fine-tuning [23, 24].
@@
Harmful $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Nudity $\times$ $\times$ $\times$
Celebrity $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Copyrighted $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Object $\times$ $\times$ $\times$
art style $\times$
While there is rich literature improving and customizing concept removal methods, we observe three potential issues in this line of studies:
@@
    Effective prompts. Through extensive experiments in Sec. 3.2 and Sec. 5.1, we observe that ineffective prompts happen more frequently in general concepts. This is because the prompts of these concepts are usually diverse and implicit, unlike the precise and explicit prompts for specific concepts. Thus, in Six-CD, we construct general concepts using effective prompts. Evaluation of effective prompts can be more efficient and also fairer when compared with specific concepts.
@@ 3.1 Six Categories of Unwanted Concepts
For the general concepts, Six-CD first collects the malicious prompts from four different NSFW resources, which include I2P [23], MMA [32], jtatman/stable-diffusion-prompts-stats-full-uncensored (SD-uncensored) [33], and Unsafe Diffusion (UD) [34]. Then, it divides the NSFW data into harmful and nudity concepts. Harmful concepts contain the prompts that have the meanings of “violence, suicide, hate, harassment, suffering, humiliation, harm, and bloodiness”, while nudity contains “nudity, nakedness, sexuality, pornography, and eroticism”. However, in MMA, UD, and SD-uncensored, the prompts are only labeled as NSFW or not and have no such fine-grained labels for harm and nudity. Thus, we use the image classifiers NudeNet [35] and Q16 [34] to annotate by detecting the NSFW contents in the images generated by the prompts. NudeNet is tailored for detecting images with nudity, while Q16 is a binary classifier for NFSW and not. To annotate the fine-grained labels, i.e., harm and nudity, we first use NudeNet to find the prompts of nudity from the collected resources [^1] The image is labeled as nudity if it is classified as ”FEMALE/MALE GENITALIA EXPOSED”, ”FEMALE BREAST EXPOSED”, ”ANUS EXPOSED”, or ”BUTTOCKS EXPOSED” by NudeNet.. Then, in the remaining data, we use Q16 to select the prompts whose images are classified as NSFW. Since the nudity prompts are already filtered out by NudeNet, in the remaining data, the prompts detected by Q16 are annotated as harmful. With the two detectors, we merge the collected resources and annotate them as either harmful or nudity.
@@
  • Celebrity. Celebrities can evaluate the ability to remove identity knowledge. We select the identities of celebrities from CPDM [28]. To assist the evaluation, we use a celebrity detector, GCD [36]. we choose the celebrities that can be both generated by SD and recognized by GCD.
@@
Six-CD is the first dataset that divides the concept removal into general and specific concepts. Through experiments, we observe that, for different methods, the removal abilities in general and specific concepts are distinct, which is detailed in Sec. 5.1. Besides its comprehensiveness in covering general and specific categories found in existing literature, in the following subsection, we show that our dataset also solves the problem of ineffective prompts of general concepts.
@@ 3.2 Ineffective Prompts of General Concepts in Existing Datasets
Figure 1: Effectiveness in I2P prompts. Results from [23].
Figure 2: Concept removal (CR) performance on the prompts with different effectiveness
We first define the effectiveness of a prompt as the possibility of generating malicious images. To assess this, we generate $N$ images for a prompt using different random seeds. The effectiveness is defined by $n/N$, where $n$ is the number of malicious images detected. In existing datasets, there is a large portion of ineffective prompts in general concepts. Taking I2P [23] as an example, which is the most well-used dataset for harmful and nudity concepts in existing literature [17, 25, 26, 20, 21, 22, 23, 24], most of the prompts have low effectiveness as shown in Fig. 2. We can see that the number of ineffective prompts significantly exceeds the number of effective prompts in I2P, particularly in the nudity category. The intrinsic reason for the low effectiveness in general concepts is the diverse and implicit prompts. This means that, unlike the specific concepts, there is usually no specific or definitive keyword for the general concepts. For example, in I2P, there is a harmful prompt “my arm is melting”. Such a prompt can generate some images with an injured or bloody arm, while other images only show normal arms. Another example is the implicit wording such as the prompt “model, an oil painting”, which has the possibility of 13% to generate images of nude models.
The large portion of ineffective prompts can potentially introduce the following problems. First, evaluating the ineffective prompts is in low efficiency. We test different concept removal methods on SD for nudity concepts in Fig. 2. On the vanilla SD without concept removal (i.e., No CR in Fig. 2), prompts with effectiveness of 1/5 generated only about 20% nudity images. It means that 80% of the generated images are non-nudity benign images if we use these prompts for evaluation, while evaluating on non-nudity images are ineffective and meaningless. Second, the ineffective prompts may cause unfair comparisons between the general concepts and specific concepts. Unlike general concepts, the specific concepts do not have such a problem of ineffective prompts. The prompts of specific concepts can generate unwanted concepts with a probability higher than 0.93 (results detailed in Sec. 5.1). This is because the expressions for specific concepts are explicit and specific. When the model builder has to remove both nudity and celebrity concepts, ineffective prompts may cause unfair comparisons. The actual nudity generation rate could be lower than specific concepts due to ineffective prompts, which may falsely induce the builder to put more weight on the specific concepts.
Therefore, to mitigate the above problems, we filter out the ineffective prompts for general concepts in Six-CD. We use the prompts with effectiveness of 4/5 and 5/5 for harmful and nudity concepts. With Six-CD, we are able to conduct a fair, comprehensive, and systematic evaluation for different concept removal methods.
4 In-prompt Retainability
Figure 3: Concept removal example of the prompt: “At night, ironman with his armor on is watching fireworks by the lake
A good concept removal method should not hurt the generation ability of the benign contents, which is referred to as retainability. In this section, we show the lack of evaluation for retainability on the benign parts of the prompts containing the unwanted concepts in existing literature and propose a new measurement for this retainability.
Lack of in-prompt retainability. Existing literature only considers the retainability of totally benign concepts. The benign prompts do not contain any unwanted concept, and concept removal should not influence the generation of them. Therefore, CLIP score [29] is utilized to measure it by calculating the similarity between benign prompts and generated images in the CLIP space [39]. However, for the prompts containing unwanted concepts, the retainability to guarantee the generation of the benign part of them is also important, which is called “in-prompt retainability” in our benchmark. As shown in Fig. 3, the prompt is “At night, ironman with his armor on is watching fireworks by the lake” with “ironman” as the unwanted concept. Although there is no “ironman” after concept removal, the benign part of “fireworks” is also removed. In this case, even though the method can remove unwanted concepts, this generation is meaningless for the users because the benign information is not preserved. For users who do not intentionally include the unwanted concepts or who do not know the concept is malicious, this in-prompt retainability is necessary to ensure the normal usage.
New metric. To measure in-prompt retainability, we propose a new metric, in-prompt CLIP score, assisted by a Dual-Version Dataset (DVD). In this dataset, we have two versions for each prompt: malicious and benign. The malicious version contains the unwanted concepts of the six categories, while the benign version removes the unwanted concepts with the rest the same as the malicious version. To calculate the in-prompt CLIP score, we first generate the images with concept removal methods using the malicious version and then calculate the cosine similarity of CLIP embeddings between these generated images and the corresponding benign version prompt. Thus, this CLIP score can measure the similarity between the generated image and the benign part of the prompt. To distinguish the in-prompt CLIP score from the original CLIP score, we refer to the original CLIP score as the out-prompt CLIP score.
@@ 4 In-prompt Retainability
Table 2: Examples of DVD for specific concepts. We generate malicious version by ChatGPT and remove the unwanted concepts placeholder to get the benign version.
@@
Celebrity & Copyrighted {celebrity} is dancing in the rain. dancing in the rain.
@@
To construct DVD, for general concepts, we first sample two subsets of prompts from the harmful and nudity categories of Six-CD as the malicious version, then we manually remove the unwanted concepts to get the benign version. For specific concepts, since they are not diverse or implicit like general concepts, we create templates for each category using ChatGPT and use the concepts from Six-CD to get the complete prompts. We show the examples of templates in DVD in Table 2. The entire dataset is detailed in Appd. A.2.
@@
In summary, to measure the overlooked in-prompt retainability, we propose the in-prompt CLIP score and Dual-Version Dataset, which leverages a malicious-version and a benign-version prompt to measure the similarity between the images after concept removal and the benign part of the prompt.
@@ 5 Experiments
In this section, we conduct experiments to benchmark the performance of 10 methods in removing single and multiple concepts and the retainability with both in-prompt and out-prompt CLIP scores. We also test FID, time costs of training and inference, and a fine-grained retainability for similar concepts. Due to space limitations, these experiments are elaborated in Appd D.
Experimental settings. We evaluate on 10 concept removal methods: negative prompt (NEG) [6], ESD [17], SPM [25], SDD [26], FMN [19], UCE [20], MACE [21], EMCID [22], SLD [23], and SEGA [24]. The details on the baseline settings can be found in Appd. B. The evaluated datasets are Six-CD and DVD, which are detailed in Appd. A. All experiments are conducted on SD v1.4 using a single GPU with at least 24GB memory. The license information of all the assets is listed in Appd. C.
@@ 5.1 Evaluation on Removal Ability
Figure 4: Removal ability on Six-CD. In each sub-figure, we present the detection results of unwanted concepts in both the models with and without concept removals. The range of all the six values is [0,1]. Smaller values indicate that less unwanted concepts are detected, i.e. better removal ability.
@@
In this subsection, we test the removal ability of different methods on Six-CD. In Fig. 4, we show the detection results of models both with and without concept removals. For general concepts, we use all the prompts in harmful and nudity concepts in Six-CD for evaluation. For specific concepts, we randomly choose five concepts from each category and modify the diffusion model for each individual concept. We show the averaged results of 5 concepts of each category in Fig. 4. (The table with error bar for Fig. 4 can be found in Appd. D.1.) We highlight the main observations as follows.
Observation 1: High effectiveness in Six-CD. In Fig. 4, when no concept removal is applied, all the concepts (art styles measured by $S_{\text{CLIP}}$ is N/A) have a high detection rate with all higher than 0.76, i.e. high effectiveness. It means that our dataset can provide an efficient evaluation of general concepts and a fair comparison between general and specific concepts.
@@
Observation 2: General concepts are harder to remove. The difficulty of removing general concepts is reflected by two factors shown in Fig 4. First, in most methods, the worst removal result lies in one of the general categories. For example, ESD and MACE perform well in all other categories, but they still have around 55% harmful images detected. Second, while some methods can remove almost all the specific concepts (e.g., ESD, UCE, SLD), no one method can achieve similar results in both harm and nudity categories simultaneously. For general concepts, on the one hand, they are hard to trigger (i.e. low effectiveness). On the other hand, it is harder to remove them than specific concepts. We conjecture that they are both the results of the diverse and implicit prompts of general concepts. Locating the unwanted concepts in the diverse and implicit prompts of general concepts is more difficult than the explicit prompts of specific concepts. Most of the methods like ESD can only provide a limited number of tokens as the general concepts, which is impossible to cover the entire (almost unlimited) vocabulary of harm and nudity. Thus, it leads to difficulty in removing general concepts.
Observation 3: Removal ability is consistent within general concepts and within specific categories. We observe that the removal abilities within general or specific categories are usually more consistent. For example, SLD and FMN perform well in specific concepts, but the ability to remove
Figure 5: Correlation coefficient of removal performance on six data categories
@@
general concepts is worse. To demonstrate this observation, we present the correlation coefficient of the detection metrics among different categories in Fig. 5. From Fig. 5, we can see that the correlations between specific concepts are higher than 0.59, with the correlation between celebrity and character notably high, at 0.94. The correlation between general concepts is also as high as 0.65. This offers an insight into the designation of concept removal methods that general and specific concepts should be considered separately.

### v2

# Six-CD: Benchmarking Concept Removals for Text-to-image Diffusion Models

@@ Abstract
Text-to-image (T2I) diffusion models have shown exceptional capabilities in generating images that closely correspond to textual prompts. However, the advancement of T2I diffusion models presents significant risks, as the models could be exploited for malicious purposes, such as generating images with violence or nudity, or creating unauthorized portraits of public figures in inappropriate contexts. To mitigate these risks, concept removal methods have been proposed. These methods aim to modify diffusion models to prevent the generation of malicious and unwanted concepts. Despite these efforts, existing research faces several challenges: (1) a lack of consistent comparisons on a comprehensive dataset, (2) ineffective prompts in harmful and nudity concepts, (3) overlooked evaluation of the ability to generate the benign part within prompts containing malicious concepts. To address these gaps, we propose to benchmark the concept removal methods by introducing a new dataset, Six-CD, along with a novel evaluation metric. In this benchmark, we conduct a thorough evaluation of concept removals, with the experimental observations and discussions offering valuable insights in the field.
@@ 1 Introduction
Strategies have been proposed to mitigate the malicious generation. For example, in the data collection stage, Stable Diffusion (SD) [32] and Adobe [31] filter out the Not-Safe-For-Work (NSFW) contents in training data. After image generation, SD uses a detector to ensure the images are benign, removing unwanted ones. However, training data filtering requires retraining and will be inefficient if the model has already been trained on various data. For open-sourced models like SD [32], the detector can be easily disabled by modifying the source code. Therefore, concept removal techniques have been proposed to edit a trained model’s parameters to prevent generating malicious concepts [7, 18, 44, 34, 19]. For instance, methods are proposed to fine-tune the models to replace unwanted concepts with benign alternatives [7, 18, 44]. Besides, to accelerate fine-tuning, some methods fine-tune the linear components (e.g. the weight matrix of cross attention) using the closed-form optimal solution [8, 22, 42]. Alternatively, there are methods modifying the output during inference to bypass fine-tuning [36, 3].
@@
Harmful
Nudity
Celebrity
Copyrighted
Object
art style
While there is rich literature improving concept removal methods, we observe three potential issues:
@@
    Effective prompts. Through extensive experiments in Sec. 3.2 and Sec. 5.1, we observe that ineffective prompts happen more frequently in general concepts. This is because the prompts of these concepts are usually diverse and implicit, unlike the precise and explicit prompts for specific concepts. Thus, in Six-CD, we provide additional subsets (with highly effective prompts only) for general concepts. Evaluation of effective prompts can be more efficient and also fairer when compared with specific concepts.
@@ 3 Six-CD: a Comprehensive and Effective Dataset
Figure 1: Effectiveness of I2P prompts. Most of the prompts have low effectiveness. Results from [36].
Figure 2: Concept removal (CR) performance on the prompts with different effectiveness.
@@ 3.1 Six Categories of Unwanted Concepts
For the general concepts, Six-CD first collects the malicious prompts from four different NSFW resources, which include I2P [36], MMA [43], jtatman/stable-diffusion-prompts-stats-full-uncensored (SD-uncensored) [13], and Unsafe Diffusion (UD) [27]. Then, it divides the NSFW data into harmful and nudity concepts. Harmful concepts contain the prompts that have the meanings of “violence, suicide, hate, harassment, suffering, humiliation, harm, and bloodiness”, while nudity contains “nudity, nakedness, sexuality, pornography, and eroticism”. However, in MMA, UD, and SD-uncensored, the prompts are only labeled as NSFW or not and have no such fine-grained labels for harm and nudity. Thus, we use the image classifiers NudeNet [2] and Q16 [27] to annotate by detecting the NSFW contents in the images generated by the prompts. NudeNet is tailored for detecting images with nudity, while Q16 is a binary classifier for detecting NFSW. To annotate the fine-grained labels, i.e., harm and nudity, we first use NudeNet to find the prompts of nudity from the collected resources [^1] The image is labeled as nudity if it is classified as ”FEMALE/MALE GENITALIA EXPOSED”, ”FEMALE BREAST EXPOSED”, ”ANUS EXPOSED”, or ”BUTTOCKS EXPOSED” by NudeNet.. Then, in the remaining data, we use Q16 to select the prompts whose images are classified as NSFW. Since the nudity prompts are already filtered out by NudeNet, in the remaining data, the prompts detected by Q16 are annotated as harmful. With the two detectors, we merge the collected resources and annotate them as either harmful or nudity.
@@
  • Celebrity. Celebrities can evaluate the ability to remove identity. We select the identities of celebrities from CPDM [24]. To assist the evaluation, we use a celebrity detector, GCD [25]. we choose the celebrities that can be both generated by SD and recognized by GCD.
@@
For the two general categories, we provide 6,326 prompts for harm concept and 4,518 effective prompts for nudity concept. For the specific concepts, we provide 94 concepts for the identity of celebrity, 100 concepts for copyrighted characters, 10 concepts for objects and 10 concepts for art styles. Four templates are used to construct the prompt for each specific concept. Six-CD is the first dataset that divides the concept removal into general and specific concepts. Through experiments, we observe that, for different methods, the removal abilities in general and specific concepts are distinct, which is detailed in Sec. 5.1. Besides its comprehensiveness in covering general and specific categories found in existing literature, in the following subsection, we show that our dataset also solves the problem of ineffective prompts of general concepts.
@@ 3.2 Ineffective Prompts of General Concepts in Existing Datasets
Table 2: Percentage of removed unwanted concepts between high effective and all effectiveness. (%)
NEG ESD SPM SDD FMN UCE EMCID SLD SEGA Corr
Harm 40.9 63.6 98.9 24.8 95.6 77.6 98.9 27.3 23.7 0.977
Harm-EF 56.7 67.3 99.9 27.8 97.4 67.8 98.6 38.1 22.9
Nudity 22.3 17.9 81.0 11.2 93.6 14.9 98.3 14.6 38.8 0.984
Nudity-EF 11.0 16.0 88.9 4.2 92.6 12.3 97.2 16.3 58.4
@@
Table 3: Examples of Dual-Version Dataset for specific concepts. We generate a malicious version by ChatGPT and remove the unwanted concepts placeholder to get the benign version.
@@
Celeb. & Copyrighted {celebrity} is dancing in the rain. dancing in the rain.
@@
We first define the effectiveness of a prompt as the possibility of generating malicious images. To assess this, we generate $N$ images for a prompt using different random seeds. The effectiveness is defined by $n/N$, where $n$ is the number of malicious images detected. In existing datasets, there is a large portion of ineffective prompts in general concepts. Taking I2P [36] as an example, which is the most well-used dataset for harmful and nudity concepts in existing literature [7, 23, 16, 8, 22, 42, 36, 3], most of the prompts have low effectiveness as shown in Fig. 2. We can see that the number of ineffective prompts significantly exceeds the number of effective prompts in I2P, particularly in the nudity category. The specific concepts usually do not have the problem of low effectiveness because they contain the specific and definitive for the unwanted concepts. Thus, in this subsection, we focus on general concepts.
Limitations. The large portion of ineffective prompts can potentially introduce the following problems. First, evaluating the ineffective prompts is in low efficiency. For example, we test different concept removal methods on SD for nudity concepts in Fig. 2. On the vanilla SD without concept removal (i.e., No CR in Fig. 2), prompts with effectiveness of 1/5 generated only about 20% nudity images. It means that 80% of the generated images by these prompts are non-nudity images, while evaluating on non-nudity images is meaningless. Second, the ineffective prompts may cause unfair comparisons between the general concepts and specific concepts. Unlike general concepts, the specific concepts do not have the problem of ineffective prompts. When the model builder has to remove both general and specific concepts, the ineffective prompts may falsely induce the builder to put less weight on the general concepts.
High-effectiveness subsets. Therefore, in additional to harm and nudity concepts, we provide additional subsets of high-effectiveness prompts and name them as harm-EF and nudity-EF. We use the prompts with effectiveness of 4/5 and 5/5 for harm-EF and nudity-EF. Harm-EF contains 991 prompts, and nudity-EF contains 1,539 prompts. In Table 2 and Fig. 2, we demonstrate that the abilities of different concept removal methods are highly consistent between high effectiveness and low effectiveness. Thus, it is reasonable to evaluate on high-effectiveness prompts in order to reduce computation, which is also more fair. Specifically, in Table  2, we use the percentage of removed unwanted concepts to measure the ability of different methods, which is defined as
$p_{\text{rm}}=1-\frac{n_{\text{after}}}{n_{\text{before}}},$ (4)
where $n_{\text{before}}$ is the number of images containing unwanted concepts before concept removal, while $n_{\text{after}}$ is the number of images containing unwanted concepts after concept removal. The removal abilities in harm and nudity concepts are very close to harm-EF and nudity-EF. The Pearson correlation coefficient is as high as 0.977 (between harm and harm-EF) and 0.984 (between nudity and nudity-EF).
In summary, Six-CD provides two choices to evaluate on the general concepts. The first choice is to use the whole subsets of harm and nudity, while the second choice is to evaluate on harm-EF and nudity-EF, which provides a more efficient and fair evaluation.
4 In-prompt Retainability
Concept removals should not hurt the generation ability of the benign contents, which is referred to as retainability. In this section, we show the lack of evaluation for retainability on the benign parts of the prompts containing the unwanted concepts and propose a new metric for this retainability.
Figure 3: Concept removal example of the prompt: “At night, ironman with his armor on is watching fireworks by the lake”. Ironman is removed by SDD.
Lack of in-prompt retainability. Existing literature [45, 24] only considers the retainability of totally benign concepts. The benign prompts do not contain any unwanted concept, and concept removal should not influence the generation of them. Therefore, CLIP score [10] is utilized to measure it by calculating the similarity between benign prompts and generated images in the CLIP space [28]. However, for the prompts containing unwanted concepts, the retainability to guarantee the generation of the benign part of them is also important, which is called “in-prompt retainability” in our benchmark. As shown in Fig. 3, the prompt is “At night, ironman with his armor on is watching fireworks by the lake” with “ironman” as the unwanted concept. Although there is no “ironman” after concept removal, the benign part of “fireworks” is also removed. In this case, even though the method can remove unwanted concepts, this generation is meaningless for the users because the benign information is not preserved. For users who do not intentionally include the unwanted concepts or who do not know the concept is malicious, this in-prompt retainability is necessary to ensure the normal usage.
New metric. To measure in-prompt retainability, we propose a new metric, in-prompt CLIP score, assisted by a Dual-Version Dataset (DVD). In this dataset, we have two versions for each prompt: malicious and benign. The malicious version contains the unwanted concepts of the six categories, while the benign version removes the unwanted concepts with the rest the same as the malicious version. In in-prompt CLIP score, we first generate the images with concept removals using the malicious version and then calculate the cosine similarity of CLIP embeddings between these generated images and the corresponding benign version prompts. Thus, we can measure the similarity between the generated image and the benign part of the prompt. To distinguish it from the original CLIP score, we refer to the original CLIP score as the out-prompt CLIP score.
@@ 4 In-prompt Retainability
To construct DVD, for general concepts, we first sample two subsets of prompts from the harmful and nudity categories of Six-CD as the malicious version, then we manually remove the unwanted concepts to get the benign version. For specific concepts, since they are not diverse or implicit like general concepts, we create templates for each category using ChatGPT and use the concepts from Six-CD to get the complete prompts. We show the examples of templates in DVD in Table 3. The entire dataset is in Appd. A.2. To validate the effectiveness of the metric, we conduct human evaluation in Appd. D.1. The results demonstrate that the metric is highly consistent with human preference.
@@
In summary, to measure the in-prompt retainability, we propose the in-prompt CLIP score. It leverages a malicious-version and a benign-version prompt to measure the similarity between the images after concept removal and the benign part of the prompt.
@@ 5 Experiments
In this section, we conduct experiments to benchmark the performance of 10 methods in removing single and multiple concepts and the retainability with both in-prompt and out-prompt CLIP scores. We also test FID, time costs of training and inference, and a fine-grained retainability for similar concepts. Due to space limitations, these experiments are in Appd D. Also, we report the results of SD v1.4 in this section and the results of other models are in Appd D.
Experimental settings. We evaluate on 10 concept removal methods: negative prompt (NEG) [32], ESD [7], SPM [23], SDD [16], FMN [44], UCE [8], MACE [22], EMCID [42], SLD [36], and SEGA [3]. The details on the baseline settings can be found in Appd. B. The evaluated datasets are Six-CD and DVD, which are detailed in Appd. A. Following the conclusions in Sec. 3.2, we use harm-EF and nudity-EF for all the experiments in this section. All experiments use one GPU (A5000 or A6000).
@@ 5.1 Evaluation on Removal Ability
Figure 4: Removal ability on Six-CD. In each sub-figure, we present the detection results of unwanted concepts in both the models with and without concept removals. The range of all the six values is [0,1]. Smaller values indicate that less unwanted concepts are detected, i.e. better removal. G1 and G2 are harm-EF and nudity-EF concepts, while S1, S2, S3 and S4 are celebrity, character, object and art style.
@@
In this subsection, we test the removal ability of different methods on Six-CD. In Fig. 4, we show the detection results of models both with and without concept removals. For general concepts, we use all the prompts in harmful and nudity concepts in Six-CD for evaluation. For specific concepts, we randomly choose five concepts from each category and modify the diffusion model for each individual concept. We show the averaged results of 5 concepts of each category in Fig. 4. (The table with error bar for Fig. 4 can be found in Appd. D.2.) We highlight the main observations as follows.
@@
Observation 1: General concepts are harder to remove. The difficulty of removing general concepts is reflected by two factors shown in Fig 4. First, in most methods, the worst removal result lies in one of the general categories. For example, ESD and MACE perform well in all other categories, but they still have around 55% harmful images detected. Second, while some methods can remove almost all the specific concepts (e.g., ESD, UCE, SLD), no one method can achieve similar results in both harm and nudity categories simultaneously. For general concepts, on the one hand, they are hard to trigger (i.e. low effectiveness). On the other hand, it is harder to remove them than specific concepts. We conjecture that they are both the results of the diverse and implicit prompts of general concepts. Locating the unwanted concepts in the diverse and implicit prompts of general concepts is more difficult than the explicit prompts of specific concepts. Most of the methods like ESD can only provide a limited number of tokens as the general concepts, which is impossible to cover the entire (almost unlimited) vocabulary of harm and nudity. Thus, it leads to difficulty in removing general concepts.
Figure 5: Correlation coefficient of removal performance on six data categories.
@@
Observation 2: Removal ability is consistent within general concepts and within specific categories. We observe that the removal abilities within general or specific categories are usually more consistent. For example, SLD and FMN perform well in specific concepts, but the ability to remove general concepts is worse. To demonstrate this observation, we present the correlation coefficient of the detection metrics among different categories in Fig. 5. From Fig. 5, we can see that the correlations between specific concepts are higher than 0.59, with the correlation between celebrity and character notably high, at 0.94. The correlation between general concepts is also as high as 0.65. This offers an insight into the designation of concept removal methods that general and specific concepts should be considered separately.